DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 21 Raitt Hall 121 3:45 PM

Fuglede’s conjecture and related problems

Izabella Laba (University of British Columbia)

A set E in R^{n} is called spectral if L^{2}(E) admits an orthogonal basis of exponential functions. Fuglede conjectured in 1974 that E is spectral if and only if it tiles R^{n} by translations. I will survey some recent work related to the conjecture.